OpenStudy (anonymous):

what is the value of x in: -x^3+5x^2-2x-4=0

OpenStudy (anonymous):

Either sign is incorrect or calculator wont work for this question, means you have to apply other methods

OpenStudy (anonymous):

In whatever your lecture teaches you

OpenStudy (anonymous):

hahaha really thanks .. will may be i got it wrong from the start the question is find the critical point of : y=(x^3-2x^2-x+2)over (x-2)

OpenStudy (anonymous):

\[y=\frac{ x ^{3}-2x ^{2} -x+2}{x-2}\]

OpenStudy (anonymous):

Simplifying y = (x3 + -2x2 + -1x + 2) Reorder the terms: y = (2 + -1x + -2x2 + x3) Remove parenthesis around (2 + -1x + -2x2 + x3) y = 2 + -1x + -2x2 + x3 Solving y = 2 + -1x + -2x2 + x3 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Simplifying y = 2 + -1x + -2x2 + x3

OpenStudy (anonymous):

y = 2 + -1x + -2x2 + x3 = [ (2-X) + x^2 (x-2) ] / (x-2) X= +2 or -2 or 0

OpenStudy (anonymous):

thanks